OpenStudy (anonymous):

Please Help! 8th Grade Pre-Algebra What is the product in lowest terms? 1/4 * 3/5

OpenStudy (anonymous):

just multiply straight across... 1/4 * 3/5 = 3/20 (which will not reduce)

OpenStudy (anonymous):

1/4*3/5=3/20 decimal form .15

OpenStudy (anonymous):

Thank You both! Great answers(:

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

:)